# Minimax Hypothesis Testing for the Bradley-Terry-Luce Model

### v1

@@ VI-B Proof of Proposition 9
$=\sum_{k=1}^{n}p_{jk}-\frac{\pi_{k}}{\pi_{j}+\pi_{k}}+\sum_{k=1}^{n}\frac{\pi_{k}}{\pi_{i}+\pi_{k}}-p_{ik}+(\pi_{i}-\pi_{j})\sum_{k=1}^{n}\frac{\pi_{k}}{(\pi_{j}+\pi_{k})(\pi_{i}+\pi_{k})}$
@@
Part 2: The construction of the matrix $P\in(0,1)^{n\times n}$ is as follows. (In the proof, we drop the subscript $n$ in the sub-sequence $P_{n}$ for brevity.) For simplicity, we assume $n$ is even; otherwise, we can replace $n/2$ in the construction with $\lceil n/2\rceil$ to get the corresponding results. Below, we define the pairwise comparison matrix $P_{n}$ for $j\geq i$. The rest of the values can be obtained by skew-symmetric condition $p_{ij}+p_{ji}=1$.
@@
where $l$ is a integer less than $n/2$ and will be defined below. We are primarily interested in the first two rows of $P$. Their respective Borda scores are $\tau_{1}(P)=\frac{n}{2}-\eta$ and $\tau_{2}(P)=\frac{n}{2}-\eta+\alpha$ and hence $\Delta\tau=\tau_{1}(P)-\tau_{2}(P)=-\alpha$. Clearly, under Borda count ranking item $2$ is ranked higher than $1$. We will show that $1$ is always ranked higher than $2$ under the BTL ranking. Moreover, we will show that the deviation of $P$ from the BTL condition decays as
$\left\|\Pi P+P\Pi-{\mathbf{1}}_{n}\pi^{\mathrm{T}}\right\|_{\mathrm{F}}\leq\frac{c}{\sqrt{n}},$
@@
Here, $e$ denotes a variable and should not be confused with the mathematical constant. The above equations and the fact that $\pi$ is a probability vector gives
@@
$a\left(\frac{n-1}{2}+\eta\right)=\frac{b}{2}+\frac{c}{2}\left(\frac{n}{2}-2\right)+d\left(\frac{n}{2}-l\right)\left(\frac{1}{2}-\frac{2\eta}{n}\right)+le\left(\frac{1}{2}-\frac{2\eta}{n}\right),$
$b\left(\frac{n-1}{2}+\eta-\alpha\right)=\frac{a}{2}+\frac{c}{2}\left(\frac{n}{2}-2\right)+\frac{d}{2}\left(\frac{n}{2}-l\right)+le\left(\frac{1}{2}-\frac{\eta-\alpha}{l}\right),$
@@
$d\left(\frac{n-1}{2}-\frac{2\eta}{n}\right)=a\left(\frac{1}{2}+\frac{2\eta}{n}\right)+\frac{b}{2}+\frac{c}{2}\left(\frac{n}{2}-2\right)+\frac{d}{2}\left(\frac{n}{2}-l-1\right)+\frac{le}{2}.$
@@
$\|\Pi P+$ $P\Pi-{\mathbf{1}}_{n}\pi^{\mathrm{T}}\|_{\mathrm{F}}=\|\frac{1}{2}(\pi{\mathbf{1}}_{n}^{\mathrm{T}}+{\mathbf{1}}_{n}\pi^{\mathrm{T}})+\Pi Q+Q\Pi-{\mathbf{1}}_{n}\pi^{\mathrm{T}}\|_{\mathrm{F}}$
$\leq\frac{1}{2}\|\pi{\mathbf{1}}_{n}^{\mathrm{T}}-{\mathbf{1}}_{n}\pi^{\mathrm{T}}\|_{\mathrm{F}}+\|\Pi Q+Q\Pi\|_{\mathrm{F}}$
$\leq\|a{\mathbf{1}}_{n}^{\mathrm{T}}-\pi^{\mathrm{T}}\|_{2}+\|b{\mathbf{1}}_{n}^{\mathrm{T}}-\pi^{\mathrm{T}}\|_{2}+\frac{1}{2}\|u{\mathbf{1}}_{n-2}^{\mathrm{T}}-{\mathbf{1}}_{n-2}u^{\mathrm{T}}\|_{\mathrm{F}}+\|\Pi Q+Q\Pi\|_{\mathrm{F}},$
@@
as $\Pi Q+Q\Pi$ has only $O(n)$ non zero terms each of which is bounded above by $2/n$. The proof then follows by a simple calculation.
To verify our calculations, we also plot the numerically calculated values in Fig. 1 which confirms the theoretical calculation that $\sqrt{n}\|\Pi P+P\Pi-{\mathbf{1}}_{n}\pi^{\mathrm{T}}\|_{\mathrm{F}}$ converges to a constant.

### v2

@@ Appendix E Stability of the BTL Assumption
$=\sum_{k=1}^{n}p_{jk}-\frac{\pi_{k}}{\pi_{j}+\pi_{k}}+\sum_{k=1}^{n}\frac{\pi_{k}}{\pi_{i}+\pi_{k}}-p_{ik}$
$\ \ \ \ \ \ \ \ \ +(\pi_{i}-\pi_{j})\sum_{k=1}^{n}\frac{\pi_{k}}{(\pi_{j}+\pi_{k})(\pi_{i}+\pi_{k})}$
@@
Part 2: The construction of the matrix $P\in(0,1)^{n\times n}$ is as follows. (In the proof, we drop the subscript $n$ in the sub-sequence $P_{n}$ for brevity.) For simplicity, we assume $n$ is even; otherwise, we can replace $n/2$ in the construction with $\lceil n/2\rceil$ to get the corresponding results. Below, we define the pairwise comparison matrix $P_{n}$ for $j\geq i$. The rest of the values can be obtained by the skew-symmetric condition $p_{ij}+p_{ji}=1$.
@@
where $l$ is a integer less than $n/2$ and will be defined below. We are primarily interested in the first two rows of $P$. Their respective Borda scores are $\tau_{1}(P)=\frac{n}{2}-\eta$ and $\tau_{2}(P)=\frac{n}{2}-\eta+\alpha$ and hence $\Delta\tau=\tau_{1}(P)-\tau_{2}(P)=-\alpha$. Clearly, under Borda count ranking, item $2$ is ranked higher than $1$. We will show that 11 is always ranked higher than $2$ under the BTL ranking. Moreover, we will show that the deviation of $P$ from the BTL condition decays as
$\left\|\Pi P+P\Pi-{\mathbf{1}}_{n}\pi^{\mathrm{T}}\right\|_{\mathrm{F}}\leq\frac{c}{\sqrt{n}}.$
@@
Here, $e$ denotes a variable and should not be confused with the mathematical constant. The above equations and the fact that $\pi$ is a probability vector give
@@
$a\left(\frac{n-1}{2}+\eta\right)=\frac{b}{2}+\frac{c}{2}\left(\frac{n}{2}-2\right)+d\left(\frac{n}{2}-l\right)\left(\frac{1}{2}-\frac{2\eta}{n}\right)$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +le\left(\frac{1}{2}-\frac{2\eta}{n}\right),$
$b\left(\frac{n-1}{2}+\eta-\alpha\right)=\frac{a}{2}+\frac{c}{2}\left(\frac{n}{2}-2\right)+\frac{d}{2}\left(\frac{n}{2}-l\right)$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +le\left(\frac{1}{2}-\frac{\eta-\alpha}{l}\right),$
@@
$d\left(\frac{n-1}{2}-\frac{2\eta}{n}\right)=a\left(\frac{1}{2}+\frac{2\eta}{n}\right)+\frac{b}{2}+\frac{c}{2}\left(\frac{n}{2}-2\right)$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\frac{d}{2}\left(\frac{n}{2}-l-1\right)+\frac{le}{2}.$
@@
$\big\|\Pi P+$ $P\Pi-{\mathbf{1}}_{n}\pi^{\mathrm{T}}\big\|_{\mathrm{F}}$
$=\left\|\frac{1}{2}(\pi{\mathbf{1}}_{n}^{\mathrm{T}}+{\mathbf{1}}_{n}\pi^{\mathrm{T}})+\Pi Q+Q\Pi-{\mathbf{1}}_{n}\pi^{\mathrm{T}}\right\|_{\mathrm{F}}$
$\leq\frac{1}{2}\left\|\pi{\mathbf{1}}_{n}^{\mathrm{T}}-{\mathbf{1}}_{n}\pi^{\mathrm{T}}\right\|_{\mathrm{F}}+\left\|\Pi Q+Q\Pi\right\|_{\mathrm{F}}$
$\leq\left\|a{\mathbf{1}}_{n}^{\mathrm{T}}-\pi^{\mathrm{T}}\right\|_{2}+\left\|b{\mathbf{1}}_{n}^{\mathrm{T}}-\pi^{\mathrm{T}}\right\|_{2}$
$\ \ \ \ \ \ \ \ \ +\frac{1}{2}\left\|u{\mathbf{1}}_{n-2}^{\mathrm{T}}-{\mathbf{1}}_{n-2}u^{\mathrm{T}}\right\|_{\mathrm{F}}+\left\|\Pi Q+Q\Pi\right\|_{\mathrm{F}},$
@@
as $\Pi Q+Q\Pi$ has only $O(n)$ non zero terms each of which is bounded above by $2/n$. The proof then follows by a simple calculation. ∎
Part 1 of Proposition 10 highlights the relationship between Borda and BTL rankings in terms of the weighted sum of entries of the error matrix. Moreover, part 2 conveys that the BTL assumption may potentially generate a “wrong” ranking (with respect to the Borda ranking) when the underlying pairwise comparison matrix is $O(1/\sqrt{n})$-separation distance away from the class of BTL models. Interestingly, this $O(1/\sqrt{n})$ deviation coincides with the critical threshold for the BTL testing problem (up to constant factors). Interestingly, this $O(1/\sqrt{n})$ deviation coincides with the critical threshold for the BTL testing problem (up to constant factors) when considering constant $k$ - we note that our stability analysis is conducted with respect to the true underlying probabilities rather than empirical probabilities which explains why our results do not exhibit dependence on $k$. It would be interesting to further explore the stability of the BTL assumption in the context of rankings in future work. We emphasize that our results here only consider stability in the complete graph case. When the induced graph is not complete, defining Borda rankings becomes ambiguous. Indeed, for a general graph, Borda rankings may not coincide with BTL rankings even under the BTL model. Finally, to verify our calculations in the proof, we also plot numerically computed values in Fig. 5 that confirm the theoretical calculation that $\sqrt{n}\|\Pi P+P\Pi-{\mathbf{1}}_{n}\pi^{\mathrm{T}}\|_{\mathrm{F}}$ converges to a constant.